x^2+y^2=1 then min value of (4-y)/(7-x)

Grade:11

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Givenx^2+y^2=1 so y=sqrt(1-x2).............(1)
Asked to find outmin value of (4-y)/(7-x)
substituting (1) in(4-y)/(7-x) we get
(4-y)/(7-x)=(4-\sqrt{1-x^{2}})/7-x
Differentiating it once and making it equal to zer we get x=-5/13,3/5
But for x=-5/13 we get second differentiation value<0
so at x=-5/13 and y=12/13 we get minimum value for above expression
So the minimum value of (4-y)/(7-x)=5/12.